### v1

@@ 6.4. Higher $\mathrm{L}^{p}$-based Sobolev regularity
    $\leq\widetilde{C}_{n,p}\left(\|\mathbf{f}\|_{0,r_{n+1};\Gamma}+\|g\|_{1,r_{n+1};\Gamma}+\|\mathbf{u}\|_{0,q_{n};\Gamma}^{2}\right).$
    We claim that this algorithm terminates, that is: there is $N\in\mathrm{N}$ such that $q_{N}\geq\frac{d}{2}$. By contradiction suppose $q_{n}<\frac{d}{2}$ for each $n\in\mathrm{N}$. Since $q_{0}>\frac{d}{3}$, it is easy to see by strong induction that $\{q_{n}\}_{n\in\mathrm{N}}$ is strictly increasing. Hence, there exists $\overline{q}>\frac{d}{3}$ such that $\lim_{n\to\infty}q_{n}=\overline{q}$. By taking limit as $n\to\infty$ in (6.37), we obtain that $\overline{q}:=\frac{d\overline{q}}{2d-3\overline{q}}$, or equivalently, $\overline{q}=\frac{d}{3}$; which is a contradiction.
    Let $N$ be the smallest integer such that $q_{N}\geq\frac{d}{2}$. From 2.6 we deduce that $\nabla_{\Gamma}\mathbf{u}\in\mathbb{L}^{d}(\Gamma)$ and $\mathbf{u}\in\mathbf{L}^{q}_{t}(\Gamma)$ for each $1<q<\infty$, whence, by Hölder’s inequality, $(\nabla_{\Gamma}\mathbf{u})\mathbf{u}\in\mathbf{L}^{q}_{t}(\Gamma)$ for each $1<q<d$. In particular, for $q=p<d$,
    $\|(\nabla_{\Gamma}\mathbf{u})\mathbf{u}\|_{0,p;\Gamma}\leq\|\nabla_{\Gamma}\mathbf{u}\|_{0,d;\Gamma}\|\mathbf{u}\|_{0,\frac{dp}{d-p};\Gamma}\leq c_{d,p}\|\mathbf{u}\|_{2,\frac{d}{2};\Gamma}^{2}\leq c_{d,p,N}\|\mathbf{u}\|_{2,r_{N};\Gamma}^{2}.$ (6.39)
    Noticing that $r_{n}=q_{n}$ for each $n\in\{1,\dotsc,N-1\}$, estimate (6.36) follows from combining (6.39) with subsequent applications of estimates (6.38) starting from n=N1n=N-1.
Step 2: Case $m\geq 1$. Assume that (6.35) holds for a fixed $m\geq 0$. Assume $\Gamma$ is of class $C^{m+4}$ and $\mathbf{f}\in\mathbf{W}^{m+1,p}_{t}(\Gamma)$ and $g\in\mathrm{W}^{m+2,p}_{\#}(\Gamma)$. By the induction hypothesis we deduce that $(\mathbf{u},\uppi)\in\mathbf{W}^{m+2,p}_{t}(\Gamma)\times\mathrm{W}^{m+1,p}_{\#}(\Gamma)$ and
$\|\mathbf{u}\|_{m+2,p;\Gamma}+\|\uppi\|_{m+1,p;\Gamma}\leq F_{m,p}(\|\mathbf{f}\|_{m,p,\Gamma},\|g\|_{m+1,p;\Gamma},\|\mathbf{u}\|_{m+1,p;\Gamma}).$
@@
To see this, we proceed via a density argument. For ease of notation herein we adopt the Einstein summation convention. We restrict ourselves to the case $\frac{d}{2}<p<d$ because the case $p\geq d$ is easier. Let $\{\mathbf{u}^{(n)}\}_{n\in\mathrm{N}}\subseteq\mathbf{C}^{m+3}_{t}(\Gamma)$ be such that $\lim_{n\to\infty}\|\mathbf{u}-\mathbf{u}^{(n)}\|_{m+2,p;\Gamma}=0$ (cf. 2.11). Since $\mathbf{u}\in\mathbf{W}^{2,p}_{t}(\Gamma)$, Sobolev embeddings $\mathbb{W}^{1,p}(\Gamma)\hookrightarrow\mathbb{L}^{\frac{dp}{d-p}}(\Gamma)$ and $\mathbf{W}^{2,p}(\Gamma)\hookrightarrow\mathbf{L}^{\infty}(\Gamma)$ (cf. 2.6 and 2.7) imply that $\lim_{n\to\infty}\|\mathbf{u}-\mathbf{u}^{(n)}\|_{0,\infty;\Gamma}=\lim_{n\to\infty}\|\nabla_{M}\mathbf{u}-\nabla_{M}\mathbf{u}^{(n)}\|_{0,\frac{dp}{d-p};\Gamma}=\|(\nabla_{\Gamma}\mathbf{u})\mathbf{u}-(\nabla_{\Gamma}\mathbf{u}^{(n)})\mathbf{u}^{(n)}\|_{0,p;\Gamma}=0$. Consequently, upon defining
$v^{(n)}_{il}$ $:=(\underline{D}_{l}P_{ij})(\underline{D}_{k}u^{(n)}_{j})u^{(n)}_{k}+P_{ij}(\underline{D}_{l}\underline{D}_{k}u^{(n)}_{j})u^{(n)}_{k}+P_{ij}(\underline{D}_{k}u^{(n)}_{j})\underline{D}_{l}u^{(n)}_{k}$
$v_{il}$ $:=(\underline{D}_{l}P_{ij})(\underline{D}_{k}u_{j})u_{k}+P_{ij}(\underline{D}_{l}\underline{D}_{k}u_{j})u_{k}+P_{ij}(\underline{D}_{k}u_{j})\underline{D}_{l}u_{k},$
it is easy to see that $\underline{D}_{l}\left[(\nabla_{\Gamma}\mathbf{u}^{(n)})\mathbf{u}^{(n)}\right]_{i}=v^{(n)}_{il}$ and
$\left\|\underline{D}_{l}\left[(\nabla_{\Gamma}\mathbf{u}^{(n)})\mathbf{u}^{(n)}\right]_{i}-v_{il}\right\|_{0,p;\Gamma}=\|v^{(n)}_{il}-v_{il}\|_{0,p;\Gamma}\xrightarrow[]{n\to\infty}0.$
With this last identity at hand, we further see that for every function $\varphi\in C^{1}(\Gamma)$ and $i,l\in\{1,\dotsc,d+1\}$, we have by integration-by-parts formula [11, Lemma B.5]
$\int_{\Gamma}\left[(\nabla_{\Gamma}\mathbf{u})\mathbf{u}\right]_{i}\underline{D}_{l}\varphi=\lim_{n\to\infty}\int_{\Gamma}\left[(\nabla_{\Gamma}\mathbf{u}^{(n)})\mathbf{u}^{(n)}\right]_{i}\underline{D}_{l}\varphi\\ =\lim_{n\to\infty}\left(-\int_{\Gamma}\underline{D}_{l}\left[(\nabla_{\Gamma}\mathbf{u}^{(n)})\mathbf{u}^{(n)}\right]_{i}\varphi+\int_{\Gamma}\operatorname*{tr}(\mathbf{B})\left[(\nabla_{\Gamma}\mathbf{u}^{(n)})\mathbf{u}^{(n)}\right]_{i}\varphi\nu_{l}\right)=-\int_{\Gamma}v_{il}\varphi+\int_{\Gamma}\operatorname*{tr}(\mathbf{B})\left[(\nabla_{\Gamma}\mathbf{u})\mathbf{u}\right]_{i}\varphi\,\nu_{l}.$
That is, $\underline{D}_{l}\left[(\nabla_{\Gamma}\mathbf{u})\mathbf{u}\right]_{i}=v_{il}$ for each $i,l\in\{1,\dotsc,d+1\}$ (cf. 2.2). From the definition of $v_{il}$ and Sobolev embeddings it is apparent that $\|v_{il}\|_{0,p;\Gamma}\lesssim\|\mathbf{u}\|_{2,p;\Gamma}^{2}$; this proves (6.40) for $m=0$. The general case $m\in\mathrm{N}$ follows by induction, using higher Sobolev embeddings.
@@
With the above claim at hand, then it is clear that $(\mathbf{u},\uppi)$ is a solution of the (weak form of) the tangent Stokes equations (4.8) with data $\overline{\mathbf{f}}:=\mathbf{f}-(\nabla_{\Gamma}\mathbf{u})\mathbf{u}\in\mathbf{W}^{m+1,p}_{t}(\Gamma)$ and $\overline{g}:=g\in\mathrm{W}^{m+2,p}_{\#}(\Gamma)$. Hence, by a direct application of 4.9 we deduce that $(\mathbf{u},\uppi)\in\mathbf{W}^{m+3,p}_{t}(\Gamma)\times\mathrm{W}^{m+2,p}_{\#}(\Gamma)$ and
$\|\mathbf{u}\|_{m+3,p;\Gamma}+\|\uppi\|_{m+2,p;\Gamma}\lesssim\|\overline{\mathbf{f}}\|_{m+1,p;\Gamma}+\|\overline{g}\|_{m+2,p;\Gamma}\leq\|\mathbf{f}\|_{m+1,p;\Gamma}+\|(\nabla_{\Gamma}\mathbf{u})\mathbf{u}\|_{m+1,p;\Gamma}+\|g\|_{m+2,p;\Gamma}\\ \lesssim\|\mathbf{f}\|_{m+1,p;\Gamma}+\|\mathbf{u}\|_{m+2,p;\Gamma}^{2}+\|g\|_{m+2,p;\Gamma}\lesssim\|\mathbf{f}\|_{m+1,p;\Gamma}+F_{m,p}(\|\mathbf{f}\|_{m,p,\Gamma},\|g\|_{m+1,p;\Gamma},\|\mathbf{u}\|_{m+1,p;\Gamma})^{2}+\|g\|_{m+2,p;\Gamma}.$
This finishes the induction step, and in consequence, the proof of 6.9.
@@
If $d=2$ and $p=\frac{4}{3}$ (hence $p^{*}=4$), then $\mathbf{W}^{1,\frac{4}{3}}(\Gamma)$ is embedded into $\mathbf{L}^{4}(\Gamma)$. Consequently, we can only ensure $(\nabla_{\Gamma}\mathbf{u})\mathbf{u}\in\mathbf{L}^{1}_{t}(\Gamma)$, which is not covered by 4.9 (higher $\mathrm{L}^{p}$-based Sobolev regularity of the tangent Stokes equations).
If $d\geq 3$ and $p=\frac{d}{2}$, then $\mathbf{W}^{1,\frac{d}{2}}(\Gamma)$ is embedded into $\mathbf{L}^{d}(\Gamma)$, and so $(\nabla_{\Gamma}\mathbf{u})\mathbf{u}\in\mathbf{L}^{q_{0}}_{t}(\Gamma)$, for $q_{0}=\frac{d}{3}$. If $d=3$, then $q_{0}=1$, which is not covered by 4.9. On the other hand, if $d\geq 4$ then the iterative process described in (missing) iii within the proof of 6.9 fails to increase the integrability of the solution $(\mathbf{u},\uppi)$. More precisely, $q_{n}=\frac{d}{3}$ for each $n\in\mathrm{N}$, whence we are left with the “suboptimal” result: $(\mathbf{u},\uppi)\in\mathbf{W}^{1,\frac{d}{3}}_{t}(\Gamma)\times\mathrm{L}^{\frac{d}{3}}_{\#}(\Gamma)$ and
@@
Should we have a stability bound of the form $\|\mathbf{u}\|_{1,p;\Gamma}\lesssim\|\mathbf{f}\|_{(\mathbf{W}^{1,p^{*}}_{t}(\Gamma))^{\prime}}+\|g\|_{0,p;\Gamma}$, then estimate (6.34) in 6.9 could be improved to
$\|\mathbf{u}\|_{m+2,p;\Gamma}+\|\uppi\|_{m+1,p;\Gamma}\leq\widetilde{F}_{m,p}(\|\mathbf{f}\|_{m,p,\Gamma},\|g\|_{m+1,p;\Gamma}).$
In particular, if $d\in\{2,3\}$, $p=2$ and $g=0$, from 6.2 (well-posedness of the Oseen equations) with $\overline{\mathbf{u}}=\mathbf{u}\in\mathbf{L}^{d}_{t,\sigma}(\Gamma)$ we know that $\|\mathbf{u}\|_{1;\Gamma}\lesssim\|\mathbf{f}\|_{(\mathbf{H}^{1}_{t}(\Gamma))^{\prime}}$ and so
$\|\mathbf{u}\|_{m+2;\Gamma}+\|\uppi\|_{m+1;\Gamma}\leq\widehat{F}_{m}(\|\mathbf{f}\|_{m,p,\Gamma}).$

### v2

@@ 5.4. Higher $\mathrm{L}^{p}$-based Sobolev regularity of tangent Navier–Stokes
    $\leq\widetilde{C}_{n,p}\left(\|\mathbf{f}\|_{0,r_{n+1};\Gamma}+\|g\|_{1,r_{n+1};\Gamma}+\|\mathbf{u}\|_{2,q_{n};\Gamma}^{2}\right).$
    Notice that if $q_{n+1}<\frac{d}{2}<p$, then $r_{n+1}=q_{n+1}$. Consequently, if there is $K\in\mathrm{N}$ such that $q_{n+1}<\frac{d}{2}$ for each $n\in\{1,\dotsc,K-1\}$, then subsequent applications of estimates (5.35) starting from n=K1n=K-1 and finishing at $n=0$, in conjunction with the simple estimates $\|\cdot\|_{0,q_{n};\Gamma}\lesssim\|\cdot\|_{0,p;\Gamma}$ and $\|\cdot\|_{1,q_{n};\Gamma}\lesssim\|\cdot\|_{1,p;\Gamma}$, yield
    $\|\mathbf{u}\|_{2,r_{K};\Gamma}\leq F_{m,p,K}\big(\|\mathbf{f}\|_{0,p;\Gamma},\|g\|_{1,p;\Gamma},\|\mathbf{u}\|_{1,p;\Gamma}\big).$ (5.36)
    We claim that the algorithm described above terminates: that is, there $N\in\mathrm{N}$ such that $q_{N}\geq\frac{d}{2}$. By contradiction suppose $q_{n}<\frac{d}{2}<p$ (hence $r_{n}=q_{n}$) for each $n\in\mathrm{N}$. Since $q_{0}>\frac{d}{3}$, it is easy to see by induction that $\{q_{n}\}_{n\in\mathrm{N}}$ is strictly increasing. Hence, there exists $\overline{q}>\frac{d}{3}$ such that $\lim_{n\to\infty}q_{n}=\overline{q}$. By taking limit as $n\to\infty$ in (5.34), we obtain that $\overline{q}:=\frac{d\overline{q}}{2d-3\overline{q}}$, or equivalently, $\overline{q}=\frac{d}{3}$; which is a contradiction.
    Let $N$ be the smallest integer such that $q_{N+1}\geq\frac{d}{2}$. Hence, for each $n\in\{0,\dotsc,N\}$ it holds that $r_{n}=q_{n}<\frac{d}{2}<p$. Since $\mathbf{u}\in\mathbf{W}^{2,q_{N+1}}(\Gamma)\subseteq\mathbf{W}^{2,\frac{d}{2}}(\Gamma)$, we deduce from 2.8 that $\nabla_{\Gamma}\mathbf{u}\in\mathbb{L}^{d}(\Gamma)$ and $\mathbf{u}\in\mathbf{L}^{q}_{t}(\Gamma)$ for each $q\in(1,\infty)$, whence, by Hölder’s inequality, $(\nabla_{\Gamma}\mathbf{u})\mathbf{u}\in\mathbf{L}^{q}_{t}(\Gamma)$ for each $q\in(1,d)$ with the estimate $\|(\nabla_{\Gamma}\mathbf{u})\mathbf{u}\|_{0,q;\Gamma}\leq C_{q}\|\mathbf{u}\|_{2,\frac{d}{2};\Gamma}^{2}$. In particular, choosing $q=p\in(1,d)$ and using (5.36) for $K=N$ let us write
    $\|(\nabla_{\Gamma}\mathbf{u})\mathbf{u}\|_{0,p;\Gamma}\lesssim\|\mathbf{u}\|_{2,\frac{d}{2};\Gamma}^{2}\leq c_{d,p,N}\|\mathbf{u}\|_{2,r_{N};\Gamma}^{2}\leq\tilde{F}_{m,p,N}\big(\|\mathbf{f}\|_{0,p;\Gamma},\|g\|_{1,p;\Gamma},\|\mathbf{u}\|_{1,p;\Gamma}\big),$ (5.37)
    which is precisely (5.33).
Step 2: Case $m\geq 1$. Assume that (5.32) holds for a fixed $m\geq 0$. Assume $\Gamma$ is of class $C^{m+3,1}$ and $\mathbf{f}\in\mathbf{W}^{m+1,p}_{t}(\Gamma)$ and $g\in\mathrm{W}^{m+2,p}_{\#}(\Gamma)$. By the induction hypothesis we deduce that $(\mathbf{u},\uppi)\in\mathbf{W}^{m+2,p}_{t}(\Gamma)\times\mathrm{W}^{m+1,p}_{\#}(\Gamma)$ and
$\|\mathbf{u}\|_{m+2,p;\Gamma}+\|\uppi\|_{m+1,p;\Gamma}\leq F_{m,p}\big(\|\mathbf{f}\|_{m,p;\Gamma},\|g\|_{m+1,p;\Gamma},\|\mathbf{u}\|_{m+1,p;\Gamma}\big).$ (5.38)
@@
To see this, we restrict ourselves to the case $p\in(\frac{d}{2},d)$ because the case $p\geq d$ is easier. First observe that, since $\mathbf{u}\in\mathbf{W}^{m+2,p}(\Gamma)$, 2.8 and 2.9 imply that $\mathbf{u}$ and all of its covariant derivatives of order less or equal than $m$ are in $\mathrm{L}^{\infty}(\Gamma)$ and that its covariant derivatives of order exactly $m+1$ belong to $\mathrm{L}^{\frac{dp}{d-p}}(\Gamma)$. Moreover, notice that $(\nabla_{\Gamma}\mathbf{u})\mathbf{u}=\mathbf{P}(\nabla_{M}\mathbf{u})\mathbf{u}$ whence, adopting Einstein summation convention, $(\nabla_{\Gamma}\mathbf{u})\mathbf{u}]_{i}=P_{ij}(\underline{D}_{k}u_{j})u_{k}$. Therefore, upon recalling 2.1 for Sobolev spaces on $\Gamma$ and the definition of differential operators on $\Gamma$ [5, Definition 3.3], we utilize the Leibniz rule for weak derivatives in Sobolev spaces in parametric domain (see, e.g. [40, Exercise 11.51(ii)]) to obtain the following explicit formula for the covariant derivatives of order $m+1$ of $(\nabla_{\Gamma}\mathbf{u})\mathbf{u}$:
$\underline{D}_{\mathbf{r}}(P_{ij}\underline{D}_{k}u_{j}u_{k})=\sum_{\begin{subarray}{c}\boldsymbol{\alpha},\boldsymbol{\beta},\boldsymbol{\gamma}\in\{0,1,\dotsc,d+1\}^{m+1}\\ \boldsymbol{\alpha}+\boldsymbol{\beta}+\boldsymbol{\gamma}=\mathbf{r}\\ \|\boldsymbol{\alpha}\|_{0}+\|\boldsymbol{\beta}\|_{0}+\|\boldsymbol{\gamma}\|_{0}=m+1\end{subarray}}\underline{D}_{\boldsymbol{\alpha}}P_{ij}\underline{D}_{\boldsymbol{\beta}}(\underline{D}_{k}u_{j})\underline{D}_{\boldsymbol{\gamma}}u_{k},\qquad\forall\mathbf{r}\in\{1,\dotsc,d+1\}^{m+1}.$
An explanation of the above compact notation follows: for any $\mathbf{t}\in\{\mathbf{r},\boldsymbol{\alpha},\boldsymbol{\beta},\boldsymbol{\gamma}\}$ we denote $\underline{D}_{\mathbf{t}}:=\underline{D}_{t_{1}}\dotsc\underline{D}_{t_{m+1}}$, with the convention that $\underline{D}_{0}$ indicates the identity operator, and $\|\mathbf{t}\|_{0}$ denotes the number of nonzero components of $\mathbf{t}$. Recalling that $p\in(\frac{d}{2},d)$ and $\Gamma$ is of class $C^{m+3,1}$, we employ Hölder’s inequalities and the aforementioned Sobolev embeddings to bound the sum above as follows. For $\boldsymbol{\beta}=\mathbf{r}$ we bound the corresponding summand as
$\|P_{ij}\underline{D}_{\mathbf{r}}(\underline{D}_{k}u_{j})u_{k}\|_{0,p;\Gamma}\lesssim\|\mathbf{u}\|_{m+2,p;\Gamma}\|\mathbf{u}\|_{0,\infty;\Gamma}\lesssim\|\mathbf{u}\|_{m+2,p;\Gamma}^{2},$
whereas for $\boldsymbol{\beta}\neq\mathbf{r}$, we instead write
$\|\underline{D}_{\boldsymbol{\alpha}}P_{ij}\underline{D}_{\boldsymbol{\beta}}(\underline{D}_{k}u_{j})\underline{D}_{\boldsymbol{\gamma}}u_{k}\|_{0,p;\Gamma}\lesssim\|\underline{D}_{\boldsymbol{\beta}}(\underline{D}_{k}u_{j})\|_{0,\frac{dp}{d-p};\Gamma}\|\underline{D}_{\boldsymbol{\gamma}}u_{k}\|_{0,d;\Gamma}\lesssim\|\mathbf{u}\|_{m+1,\frac{dp}{d-p};\Gamma}\|\mathbf{u}\|_{m+1,d;\Gamma}\lesssim\|\mathbf{u}\|_{m+2,p;\Gamma}^{2},$
because $\frac{dp}{d-p}>p$. This finishes the proof of claim (5.39).
@@
With (5.39) at hand, then it is clear that $(\mathbf{u},\uppi)$ is a solution of the (weak form of) the tangent Stokes equations (4.1) with data $\overline{\mathbf{f}}:=\mathbf{f}-(\nabla_{\Gamma}\mathbf{u})\mathbf{u}\in\mathbf{W}^{m+1,p}_{t}(\Gamma)$ and $g\in\mathrm{W}^{m+2,p}_{\#}(\Gamma)$ satisfying $\|\overline{\mathbf{f}}\|_{m+1,p;\Gamma}\lesssim\|\mathbf{f}\|_{m+1,p;\Gamma}+\|\mathbf{u}\|_{m+2,p;\Gamma}^{2}$. Hence, by a direct application of 4.8 we deduce that $(\mathbf{u},\uppi)\in\mathbf{W}^{m+3,p}_{t}(\Gamma)\times\mathrm{W}^{m+2,p}_{\#}(\Gamma)$ and, in conjunction with the estimate (5.38), that
[…]
This finishes the induction step and the proof of 5.9. ∎
@@
  • If $d=2$ and $p=\frac{4}{3}$ (hence $p^{*}=4$), then $\mathbf{W}^{1,\frac{4}{3}}(\Gamma)$ is embedded into $\mathbf{L}^{4}(\Gamma)$. Consequently, we can only ensure $(\nabla_{\Gamma}\mathbf{u})\mathbf{u}\in\mathbf{L}^{1}_{t}(\Gamma)$, which is not covered by 4.8 (higher $\mathrm{L}^{p}$-based Sobolev regularity of the tangent Stokes equations).
  • If $d\geq 3$ and $p=\frac{d}{2}$, then $\mathbf{W}^{1,\frac{d}{2}}(\Gamma)$ is embedded into $\mathbf{L}^{d}(\Gamma)$, and so $(\nabla_{\Gamma}\mathbf{u})\mathbf{u}\in\mathbf{L}^{q_{0}}_{t}(\Gamma)$, for $q_{0}=\frac{d}{3}$. If $d=3$, then $q_{0}=1$, which is not covered by 4.8. On the other hand, if $d\geq 4$ then the iterative process described in () iii within the proof of 5.9 fails to increase the integrability of the solution $(\mathbf{u},\uppi)$. More precisely, $q_{n}=\frac{d}{3}$ for each $n\in\mathrm{N}$, whence we are left with the “suboptimal” result: $(\mathbf{u},\uppi)\in\mathbf{W}^{2,\frac{d}{3}}_{t}(\Gamma)\times\mathrm{W}^{1,\frac{d}{3}}_{\#}(\Gamma)$ and
@@
Should we have a stability bound of the form $\|\mathbf{u}\|_{1,p;\Gamma}\lesssim\|\mathbf{f}\|_{(\mathbf{W}^{1,p^{*}}_{t}(\Gamma))^{\prime}}+\|g\|_{0,p;\Gamma}$, then estimate (5.31) in 5.9 could be improved to
$\|\mathbf{u}\|_{m+2,p;\Gamma}+\|\uppi\|_{m+1,p;\Gamma}\leq\widetilde{F}_{m,p}(\|\mathbf{f}\|_{m,p;\Gamma},\|g\|_{m+1,p;\Gamma}).$
In particular, if $d\in\{2,3\}$, $p=2$ and $g=0$, from 5.2 (well-posedness of the Oseen equations) with $\overline{\mathbf{u}}=\mathbf{u}$ we know that $\|\mathbf{u}\|_{1;\Gamma}\lesssim\|\mathbf{f}\|_{(\mathbf{H}^{1}_{t}(\Gamma))^{\prime}}$ and so
$\|\mathbf{u}\|_{m+2;\Gamma}+\|\uppi\|_{m+1;\Gamma}\leq\widehat{F}_{m}\big(\|\mathbf{f}\|_{m,p;\Gamma}\big).$